### v1

# Is this chart lying to me? Automating the detection of misleading visualizations

Abstract
Misleading visualizations are a potent driver of misinformation on social media and the web. By violating chart design principles, they distort data and lead readers to draw inaccurate conclusions. Prior work has shown that both humans and multimodal large language models (MLLMs) are frequently deceived by such visualizations. Automatically detecting misleading visualizations and identifying the specific design rules they violate could help protect readers and reduce the spread of misinformation. However, the training and evaluation of AI models has been limited by the absence of large, diverse, and openly available datasets. In this work, we introduce Misviz, a benchmark of 2,604 real-world visualizations annotated with 12 types of misleaders. To support model training, we also release Misviz-synth, a synthetic dataset of 81,814 visualizations generated using Matplotlib and based on real-world data tables. We perform a comprehensive evaluation on both datasets using state-of-the-art MLLMs, rule-based systems, and fine-tuned classifiers. Our results reveal that the task remains highly challenging. We release Misviz, Misviz-synth, and the accompanying code.111github.com/UKPLab/arxiv2025-misviz
Is this chart lying to me?
Automating the detection of misleading visualizations
Jonathan Tongletthanks: These authors contributed equally to this work.1,2,3, Jan Zimny11footnotemark: 11,2, Tinne Tuytelaars2, Iryna Gurevych1 1 Ubiquitous Knowledge Processing Lab (UKP Lab), Department of Computer Science, TU Darmstadt and National Research Center for Applied Cybersecurity ATHENE 2 Department of Electrical Engineering, KU Leuven 3 Department of Computer Science, KU Leuven www.ukp.tu-darmstadt.de
1 Introduction
Misleading visualizations are charts that distort the underlying data, typically by violating design principles, leading readers to draw inaccurate conclusions (Tufte and Graves-Morris, 1983; Pandey et al., 2015; Lauer and O’Brien, 2020; McNutt et al., 2020; Lo et al., 2022; Lisnic et al., 2023; Lan and Liu, 2025). While many arise from unintentional design errors, misleading visualizations are also deliberately crafted by malicious actors to spread disinformation and manipulate public understanding, especially during crises such as the COVID-19 pandemic, where misleading charts circulated widely on social media (Correll and Heer, 2017; Tartaglione and de Wit, 2025; Lisnic et al., 2023). Prior work has shown that both humans (Pandey et al., 2014, 2015; O’Brien and Lauer, 2018; Yang et al., 2021; Ge et al., 2023; Rho et al., 2023) and MLLMs (Bendeck and Stasko, 2025; Chen et al., 2025; Pandey and Ottley, 2025; Tonglet et al., 2025) are deceived by such visualizations in question-answering tasks.
The deceptive features in these charts, or misleaders (Lisnic et al., 2023), often reside in subtle details easily missed by readers, such as axis tick intervals. Furthermore, misleaders are highly diverse: the latest taxonomies identify over 70 distinct types spanning a wide variety of chart types, including bar charts, pie charts, and maps (Lo et al., 2022; Lan and Liu, 2025). In some cases, more than one misleader affects the same visualization (Lo et al., 2022). Figure 1 shows 12 real-world examples.
Figure 1: Examples of the 12 types of misleaders included in Misviz. Appendix A explains how these visualizations misrepresent their underlying data table.
Automatically classifying whether a visualization is misleading, and identifying which misleaders affect it, if any, can enable timely warnings to chart designers and readers and help prevent the spread of misinformation. This task is framed as a multi-label classification problem. While early work relied on rule-based systems called linters (Hopkins et al., 2020; Fan et al., 2022), recent studies have explored the use of MLLMs (Lo and Qu, 2025; Alexander et al., 2024). However, these approaches were evaluated on distinct datasets, which are either small or closed-access, limiting comparability and hindering progress.
In this work, we introduce Misviz, a large, diverse, and open benchmark comprising 2,604 real-world visualizations spanning 12 types of misleaders. It reflects scenarios in which vulnerable readers encounter misleading visualizations online. In Misviz, 70% of the visualizations contain up to three misleaders, while the remainder are non-misleading. To support model training, we also release Misviz-synth, a synthetic dataset generated with Matplotlib (Hunter, 2007) using real-world data tables. Misviz-synth simulates scenarios where chart designers seek assistance in identifying misleaders unintentionally introduced in their visualizations. It includes not only the visualizations but also their data tables and axis metadata, enabling the training of chart de-rendering models.
We conduct extensive experiments with three approaches: (a) state-of-the-art MLLMs, (b) a new rule-based linter that inspect axis metadata for design rule violations, and (c) new classifiers that take the visualization alone or in combination with the axis metadata as input. For (b) and (c), we fine-tune DePlot (Liu et al., 2023) to extract axis metadata. Our experiments address the following research questions (RQs). RQ1: Which type of model performs best on real-world or synthetic instances? RQ2: Can models trained on synthetic instances generalize to real-world cases? RQ3: Can models fine-tuned for axis extraction on synthetic instances generalize to real-world cases?
Our results show that MLLMs perform best on real-world visualizations, while linters and fine-tuned classifiers outperform them on synthetic ones, benefiting from the availability of training data for both axis extraction and misleader detection. While the fine-tuned DePlot can extract axes from Misviz-synth, it does not generalize well to Misviz, reducing the performance of the linter and classifier.
In summary, our contributions are as follows: (1) we introduce Misviz and Misviz-synth, the first large scale open datasets for misleading visualization detection, (2) we propose a new linter and a new classification method that combines image and extracted axis metadata as input, (3) we conduct comprehensive evaluation and error analysis.
2 Related work
Dataset Instances Misleader types Chart types % non-misleading Open access Real-world Multi-label Axis, table
MISCHA-QA (Arif et al., 2024) 8,201 4 3 49
DCDM (Maciborski et al., 2025) 24,480 5 3 51
Alexander et al. (2024) - design misleaders 1,460 7 > 5 50
Lo and Qu (2025) 150 21 > 5 16
Misvisfix (Das and Mueller, 2025) 450 74 > 5 20
Misviz-synth (ours) 81,814 11 3 22
Misviz (ours) 2,604 12 > 5 31
Table 1: Existing datasets for misleading visualization detection.
The first attempts to detect misleading visualizations relied on rule-based systems called linters (McNutt and Kindlmann, 2018; Hopkins et al., 2020; Chen et al., 2022). These linters assume the availability of the underlying data table or the chart code, which restricts their applicability to real-world scenarios. Fan et al. (2022) and Biselli et al. (2025) overcome these limitations by extracting tables with OCR tools from real-world visualizations before applying rule checks. However, the accuracy of these real-world linters depends heavily on the quality of the intermediate OCR step, which can vary widely (Biselli et al., 2025). Real-world linters have previously been evaluated in small-scale user studies, with human-in-the-loop to correct OCR errors.
Others have explored the potential of MLLMs for the task. Lo and Qu (2025) evaluated four MLLMs with different prompts on a dataset of 150 real-world visualizations, sourced from the corpus of Lo et al. (2022) for misleading cases. They found that detection accuracy decreased as more misleader types were included in the prompt. Alexander et al. (2024) focused on GPT-4 (OpenAI, 2023), using visualizations from the social media platform X (Lisnic et al., 2023). However, access to this dataset requires a paid API, and reproducibility is further hindered by the frequent removal of the posts from the platform. In a parallel work, Das and Mueller (2025) proposed a prompt with which SOTA MLLMs achieve high accuracy on a subset of the corpus of Lo et al. (2022).
Recently, Maciborski et al. (2025) fine-tuned a convolutional neural network for the task, achieving high accuracy on synthetic instances.
Table 1 compares prior datasets with Misviz and Misviz-synth. Misviz is over fifteen times larger than the dataset of Lo and Qu (2025). Unlike Alexander et al. (2024), it does not rely on paid APIs and ensures long-term access to all instances by archiving them on the Wayback Machine.222web.archive.org Misviz-synth is three to ten times larger than other synthetic datasets, and include several more misleader types (Arif et al., 2024; Maciborski et al., 2025). In contrast to other synthetic datasets, Misviz-synth provides axis metadata, which is necessary to fine-tune DePlot for axis extraction and answer our research questions.
3 Misviz
3.1 Selected misleaders
Misviz covers 12 types of misleaders, selected from the 74 categories defined in the taxonomy of Lo et al. (2022), based on three key criteria. First, we excluded misleaders that are rarely observed in real-world scenarios. To determine this, we used misleader frequency statistics from the corpus of Lo et al. (2022) and discarded all categories with fewer than three instances. Second, we removed misleaders that are more likely to confuse than deceive the reader. As noted by Lo et al. (2022), the taxonomy includes both misleaders that distort the underlying data, the focus of this work, and others that may hinder readability or clarity without altering the interpretation of the data, such as missing titles or overplotting. Third, we excluded misleaders that require specific domain knowledge to identify. For example, using red to represent Democrats and blue for Republicans in a U.S. politics chart constitutes a violation of color conventions, but detecting this misleader requires familiarity with U.S. politics. Such misleaders requires domain expertise that is not achievable with crowdworkers.
We define each selected misleader briefly below (Lo et al., 2022; Ge et al., 2023; Lan and Liu, 2025). Each of them is represented with an example in Figure 1 and in Appendix A.
Misrepresentation: the value labels displayed do not match the size of their visual encodings, e.g., bars may be drawn disproportionate to the corresponding numerical value.
3D: the visualization includes 3D effects, distorting the size of visual encodings.
Truncated axis: an axis does not start from zero, thus exaggerating differences between values.
Inappropriate use of pie chart: a pie chart does not display data in a part-to-whole relationship.
Inconsistent tick intervals: the ticks in one axis are evenly spaced but their values are not, e.g., the tick values sequence is 10, 20, 40, 45.
Dual axis: there are two independent and parallel numerical axes with different scales.
Inconsistent binning size: a variable, such as years or ages, is grouped in unevenly sized bins.
Discretized continuous variable: a continuous variable is cut into discrete categories, thus exaggerating the difference between boundary cases.
Inappropriate use of line chart: a line chart is used in unusual ways, e.g., with categorical data.
Inappropriate item order: the tick labels of an axis are sorted in an unconventional way, e.g., dates are not shown chronologically.
Inverted axis: an axis is displayed in a direction opposite to conventions.
Inappropriate axis range: the axis range is either too broad or too narrow, minimizing or exaggerating the real trend.
3.2 Data collection
We obtain visualizations from three sources.
(a) We collect instances from the corpus of Lo et al. (2022) that contain at least one of the 12 selected misleaders. We apply perceptual hashing to remove duplicates. We then manually discard instances with low resolution, those that display the name of the misleader, or those that show both a misleading chart and a corrected version side by side. Appendix B shows removed examples.
(b) We use the misleading visualizations from the website WTF Visualizations,333viz.wtf previously annotated by Lan and Liu (2025) for taxonomy construction. We align their misleader categories with those of Lo et al. (2022), as explained in Appendix C, retain only those matching one of our 12 target misleaders, and remove duplicates.
(c) We access a large collection of unlabeled visualizations from r/dataisugly and r/dataisbeautiful.444kaggle.com/datasets/bcruise/reddit-data-is-beautiful-and-ugly The former is an online community focused on sharing misleading visualizations, while the latter features high-quality examples and serves as our source of non-misleading instances. We begin by removing duplicates, then hire annotators to assign labels, as detailed in Section 3.3.
3.3 Data labeling
We split the labeling of visualizations from r/dataisugly and r/dataisbeautiful into three annotation tasks. For each task, we hired three crowdworkers from Prolific and paid them £9 per hour. To ensure sufficient familiarity with visualizations, we required annotators to hold a PhD degree. Inter-annotator agreement (IAA) was evaluated with Fleiss $\kappa$ (Fleiss, 1971) on an overlapping set of 30 instances between all crowdworkers. Appendix D shows the annotation interface and instructions.
In the first task, crowdworkers assigned one or more chart types to each visualization. Five categories were available: bar charts, line charts, maps, scatterplots, and others. The “others” category includes less frequent chart types such as treemaps. The crowdworkers achieved a high IAA (0.71).
In the second task, crowdworkers labeled visualizations from r/dataisugly with zero to three misleaders. They received detailed guidelines with definitions and examples for each misleader. Some visualizations were assigned no label and removed from the corpus, typically because they contained a misleader outside of our selected set. A moderate IAA of 0.53 is achieved. This is a reasonable outcome, given the task’s complexity with 12 categories and multiple possible combinations.
In the final task, crowdworkers verified that visualizations from r/dataisbeautiful were not misleading. Visualizations identified as containing one or more misleaders were removed from the corpus. The Fleiss $\kappa$ was 0.78, reflecting high IAA.
3.4 Data statistics
Misviz is divided into a few-shot development set (5%), a validation set for hyperparameter and prompt tuning (15%), and a held-out test set (80%). 78% of the visualizations contain one of the three main chart types: bar chart, line chart, or pie chart. While 94% of the visualizations contain a single chart type, 5 and 1% contain two or three different chart types, respectively. The split is stratified to ensure a balanced distribution of misleaders and chart types. Among the misleading visualizations, 85, 14, and 1% contain one, two, or three misleaders, respectively. The most frequent misleader is misrepresentation, present in 32% of the visualizations. This aligns with prior findings (Lo et al., 2022; Lan and Liu, 2025). The next most frequent categories are 3D effects and truncated axes, appearing in 14 and 9% of the visualizations, respectively. All other misleaders occur in 1 to 5% of the instances. Appendix E provides detailed distributions of misleaders and chart types.
Although we do not provide language distributions, it is worth noting that the visualizations span many languages. The types of images also vary significantly. While most are screenshots, some are pictures of visualizations printed on paper or displayed on screens. Appendix F provides cases of non-English and non-screenshot visualizations.
4 Misviz-synth
4.1 Data collection
We collect real-world data tables from two primary sources: Wikipedia and Our World in Data (OWID).555ourworldindata.org For Wikipedia, we use the tables included in the TabFact question answering dataset (Wenhu et al., 2020). OWID is an open-access data platform that covers a wide range of domains, including health, education, and the economy.
4.2 Synthetic visualizations generation
Figure 2: The two step process to create the synthetic visualizations of Misviz-synth based on real-world data.
We design a two-step rule-based system to generate synthetic visualizations, illustrated in Figure 2.
In the first step, given a raw real-world data table, we assign one or more types to each column. There are 10 different column types, explained in Appendix G. Based on these types, we select valid column combinations, requiring at least one numerical column and a natural key for each row. This natural key can consist of one or a combination of categorical or temporal columns. If the natural key involves multiple columns, we fix all but one to a constant value. For instance, if the numerical column is Win ratio and the natural key includes Football club and Year, we set Year to 2021 to compare the win ratios of football clubs for that year. Chart titles are automatically generated using templates based on the selected columns. We then determine which chart types, among bar, line, and pie charts, are suitable for the data. For example, line charts are only compatible with temporal data, not categorical data. Each valid configuration produces a “prepared” data table.
In the second step, we generate visualizations using hand-crafted plotting functions for each chart type–misleader pair. For each prepared data table, we identify the misleaders applicable given the column types and selected chart types. One or more misleading visualizations are generated, each with exactly one misleader from the applicable set. Additionally, non-misleading visualizations are generated for a random subset of the prepared tables.
For bar and line charts, we automatically extract axis metadata. Since pie charts do not have axes, they are excluded from this step. Axis metadata is stored as a table with four columns. (1) Seq indicates the sequential order of the tick marks along an axis. Horizontal axes are read from left to right, while vertical axes are read from bottom to top. (2) Axis indicates the name of the axis, e.g., x, y1, or y2, in case of dual axis. (3) Label is the tick label. (4) Relative position is a normalized float indicating the tick mark’s position, with spacing expressed relative to the distance between the first two tick marks. Examples of axis metadata are provided in Appendix H.
To increase visual diversity, we apply random design variations such as font size, background color, and axis label position. The complete list of variations is given in Appendix I.
Each instance is saved along with its title, data table, and axis metadata if available.
4.3 Data statistics
Misviz-synth contains 81,184 visualizations, distributed into stratified train (93%), validation (2%), and test sets (5%). Misviz-synth focuses on the three most common chart types in Misviz: bar (46%), line (47%), and pie charts (7%). Due to Matplotlib’s limitations, it does not include maps, and therefore omits the discretized continuous variable misleader, which is specific to maps. 78% of the visualizations are misleading. Unlike Misviz, each visualization in Misviz-synth contains at most one chart type and one misleader, making it slightly less diverse than Misviz.
5 Experiments
5.1 Baselines
We consider three categories of baseline models, illustrated in Figure 3.
Zero-shot MLLMs We evaluate the zero-shot capabilities on one run of the following commercial and open-weight MLLMs: GPT-4.1 and GPT-o3 (OpenAI, 2023), Gemini-1.5-Flash and Gemini-1.5-Pro (Gemini-Team, 2024), Qwen-2.5-VL in 7B, 32B, and 72B variants (Qwen-Team, 2025), and InternVL3 in 8B, 38B, and 78B variants (Zhu et al., 2025). These models were selected due to their strong performance on the ChartQA benchmark (Masry et al., 2022). The prompt includes the task description as well as the definitions of the misleaders. The prompt is provided in Appendix J.
Rule-based linter We design a linter that detects misleaders by applying rule-based checks to the axis metadata of a visualization. Each misleader corresponds to a specific rule. If a visualization passes all checks, it is classified as having no misleaders. We rely on axis metadata rather than the underlying data table because it enables detection of a broader range of misleaders: truncated axis, inverted axis, dual axis, inconsistent tick intervals, inconsistent binning size, and inappropriate item order. Only the latter two misleaders could also be detected from the data table alone. Detailed descriptions of all rule checks are provided in Appendix K. Misleaders that require visual interpretation or contextual knowledge, such as misrepresentation or inappropriate axis ranges, are not covered by the linter. We evaluate its performance using both ground truth and predicted axis metadata.
Fine-tuned classifiers We train two classifiers: one that takes only the visualization image as input, and another that combines the image with its axis metadata. Visualizations are encoded using a frozen image encoder, while axis metadata is embedded using a frozen table encoder. We use the image encoder of TinyChart, a specialized chart understanding MLLM (Zhang et al., 2024), while the table encoder is TaPas (Herzig et al., 2020). For the second classifier, the resulting image embeddings are concatenated with the [CLS] token from the axis metadata embeddings. The image embedding or concatenated embeddings are passed through a trained multi-layer perceptron (MLP) classification head. The classifiers are trained on Misviz-synth to predict one misleader per visualization, or none.
Axis extraction The linter and the second classifier require axis metadata as input. While ground truth metadata is available for Misviz-synth, this is not the case for Misviz or real-world scenarios where only the visualization image is accessible. To address this, we implement an intermediate axis extraction step using DePlot (Liu et al., 2023). Since DePlot was originally only trained for chart-to-table extraction, we fine-tune it on (image, axis metadata) pairs from Misviz-synth.
Figure 3: The three type of baselines included in the experiments. The linter and one classifier require axis extraction as an intermediate step.
5.2 Evaluation metrics
We evaluate model performance using six metrics. The first four assess the binary classification of visualizations as misleading or not: Accuracy (Acc), Precision (Pre) and Recall (Rec) for the misleading class, and macro-F1 score (F1). The remaining two metrics are computed on the subset of misleading visualizations and evaluate the correct identification of specific misleaders. Exact Match (EM) assigns a score of 1 if the set of predicted misleaders exactly matches the ground truth. Partial Match (PM) assigns a score of 1 if the predicted misleaders are a subset of the ground truth.
5.3 Implementation details
Open weight models are accessed with Transformers (Wolf et al., 2020), while commercial models are accessed via their official API. For all MLLMs, we set the temperature to 0, except for GPT-o3 where it is set to the default value of 1. Results for the fine-tuned classifiers are reported as averages over three different seeds (123, 456, 789), with standard deviations reported in Appendix L. Additional information and hyperparameter values are provided in Appendix M.
5.4 Results
We structure the results analysis in two parts. Table 2 presents a breakdown of zero-shot MLLM performance on the full test set of Misviz. In contrast, Table 3 compares the performance of all models on (1) the full test set of Misviz-synth, and (2) the subset of Misviz test instances that only include bar, line, or pie charts. This restriction ensures a fairer comparison with the linter and the fine-tuned classifiers, which only cover two and three of these chart types, respectively.
We organize our insights around the following RQs. RQ1: Which type of model performs best on real-world or synthetic instances? RQ2: Can models trained on synthetic instances generalize to real-world cases? RQ3: Can models fine-tuned for axis extraction on synthetic instances generalize to real-world cases?
Table 2 shows substantial performance differences across MLLMs, with GPT models achieving the best scores, in particular for EM and PM. We attribute this to their strong OCR capabilities. Table 3 shows that MLLMs outperform linters and fine-tuned classifiers on real-world visualizations. However, this is not the case for Misviz-synth, where they are clearly outperformed by the linter and the classifiers in terms of F1 and EM. We attribute this to two factors: (1) Misviz-synth instances often contain many axis ticks, making the detailed extraction of axis metadata very important, and (2) the axis extractor and classifiers benefit from being trained on synthetic visualizations.
Insight 1. MLLMs perform better on Misviz, while the linter and the classifiers are the best on Misviz-synth (RQ1).
Acc Pre Rec F1 EM PM
Qwen2.5-VL-7B 41.9 66.1 33.5 41.8 5.2 9.0
Qwen2.5-VL-32B 73.7 73.9 95.9 59.4 4.8 6.3
Qwen2.5-VL-72B 72.3 76.6 86.5 64.1 26.5 35.0
InternVL3-8B 63.1 68.5 86.8 45.1 22.3 28.9
InternVL3-38B 59.2 76.5 59.4 56.8 28.5 37.3
InternVL3-78B 55.6 68.1 67.8 48.0 23.8 30.9
Gemini-1.5-Flash 53.0 69.5 57.3 49.4 21.8 29.2
Gemini-1.5-Pro 75.2 84.6 78.4 72.0 35.1 40.9
GPT-4.1 84.1 84.5 94.3 79.6 53.6 64.0
GPT-o3 83.5 86.7 90.1 80.0 58.9 67.5
Table 2: Performance (%) of MLLMs on the full test set of Misviz. The best results are marked in bold.
Misviz Misviz-synth
Acc Pre Rec F1 EM PM Acc Pre Rec F1 EM PM
Zero-shot MLLMs
Qwen2.5-VL-7B 41.8 69.5 34.6 41.4 4.8 8.0 64.0 82.0 68.9 55.8 7.0 7.0
Qwen2.5-VL-32B 75.0 75.9 95.7 58.3 3.7 4.5 75.3 79.8 91.4 55.1 3.0 3.0
Qwen2.5-VL-72B 75.1 79.6 88.2 65.8 29.0 37.1 72.7 83.0 81.7 61.1 21.2 21.2
InternVL3-8B 66.1 71.3 88.6 45.1 21.7 27.4 76.7 78.5 96.6 48.9 7.5 7.5
InternVL3-38B 58.4 79.0 57.8 55.7 26.4 34.2 52.9 88.6 45.4 51.4 21.5 21.5
InternVL3-78B 55.6 70.0 67.4 46.2 23.1 29.3 62.0 78.9 68.6 53.4 23.5 23.5
Gemini-1.5-Flash 55.1 74.4 58.7 51.4 22.8 29.5 68.9 81.8 77.4 57.7 7.8 7.8
Gemini-1.5-Pro 74.2 86.4 76.6 70.5 30.7 35.5 61.3 88.6 57.8 57.9 25.1 25.1
GPT-4.1 85.2 86.0 95.0 79.7 52.3 61.8 73.4 83.6 82.0 62.3 42.0 42.0
GPT-o3 84.9 88.8 90.4 80.9 59.4 67.9 71.9 89.9 71.9 66.5 51.3 51.3
Rule-based linter
Linter w. axis (ground truth) - - - - - - 77.5 99.7 71.6 74.9 69.2 69.2
Linter w. axis (DePlot[Uncaptioned image]) 62.9 76.9 70.3 56.0 3.1 4.2 79.3 96.5 76.2 75.5 67.4 67.4
Fine-tuned classifiers
Image 69.6 72.2 94.3 45.6 27.1 34.0 79.2 83.8 90.9 66.0 76.8 76.8
Image w. axis (DePlot[Uncaptioned image]) 64.3 71.4 84.4 46.7 22.1 27.7 81.6 86.6 90.4 71.7 80.4 80.4
Table 3: Performance (%) on the test sets of Misviz and Misviz-synth for instances with bar, line, or pie charts. The best results, excluding those with ground truth axis metadata, are marked in bold.
The linter with ground truth axis metadata achieves high precision on Misviz-synth, showing the usefulness of axis metadata for misleader detection. However, the linter only covers six misleaders, negatively affecting recall and EM. Using predicted axis metadata instead of ground truth axis metadata has only a small negative effect on Misviz-synth, indicating that the fine-tuned DePlot reliably extracts axis information from Matplotlib-generated charts. However, on Misviz, the EM and PM are very low. This is because DePlot fails to generalize to real-world visualizations.
Insight 2. The linter has high precision for a subset of misleaders, but is very sensitive to axis extraction errors (RQs 1 and 3).
Insight 3. The axis extractor fine-tuned on synthetic visualizations has limited generalizability to real-world ones (RQ 3).
On Misviz-synth, the fine-tuned classifiers achieve higher EM and PM than all other baselines. The classifier with axis metadata is the strongest, further highlighting the usefulness of that input modality. Both classifiers generalize to Misviz to some extent. In that case, the classifier with axis metadata performs worse, likely due to error propagation from the axis extraction step. While the classifiers do not outperform the strongest zero-shot MLLMs on Misviz, they achieve better F1 and EM than smaller MLLMs like InternVL3-8B.
Insight 4. Classifiers trained on Misviz-synth can generalize to Misviz (RQ 2).
The best EM scores remain low on both datasets, underscoring the difficulty of identifying which misleaders affect a visualization. For binary classification, most baselines tend to favor recall over precision, with the notable exception of the linter. Future work could aim for a better balance to improve overall F1 performance.
Based on these observations, rule-based linters and fine-tuned classifiers are best suited for controlled environments where table and axis metadata are available. They can assist chart designers by automatically detecting misleaders introduced unintentionally. In contrast, for flagging misleading visualizations encountered online, where only the image is available, MLLMs are the better option.
Appendix N reports the MLLM results per data source of misleading visualization on Misviz.
5.5 Manual error analysis
Misviz Misviz-synth
Incorrect use of misleader definition 14 13
Incorrect measurement of object size 10 3
Incorrect extraction of axis metadata 2 14
Complex chart type 2 0
No detection of 3D effects 2 0
Table 4: Manual analysis of 30 prediction errors with GPT-4.1 on Misviz and Misviz-synth test sets.
We manually analyze two random samples of 30 incorrect predictions made on both datasets by GPT-4.1. Errors are categorized into different types, as reported in Table 4, with examples in Appendix O.
Around half the errors on both datasets are due to incorrect uses of misleader definitions. For instance, several Misviz-synth instances are incorrectly classified as involving a dual axis because the chart displays several lines. However, this misleader applies only when there are two distinct vertical axes. On Misviz, several instances are misclassified as inappropriate use of pie chart because the values shown do not sum to 100. GPT-4.1 fails to see that the pie slice labels are absolute values rather than percentages.
Many errors in Misviz are due to incorrect measurement of the sizes of visual encodings, such as pie slices and bar heights, which is critical to detect misrepresentation, by comparing the measured sizes with value labels to identify inconsistencies.
In Misviz-synth, many errors arise from incomplete or incorrect parsing of axis metadata. For example, GPT-4.1 frequently overlooks inconsistent tick intervals on long temporal or numerical axes with numerous tick marks.
Two additional error categories are unique to Misviz. Some real-world visualizations are complex in their structure, sometimes containing more than four charts, which overwhelms the model with excessive visual information. In other cases, GPT-4.1 fails to detect 3D effects, the only misleader that does not require to analyze any displayed labels.
6 Conclusion
In this work, we introduce two datasets for misleading visualization detection. Misviz is a real-world benchmark covering 12 types of misleaders. Misviz-synth is a synthetic dataset generated from real-world tables, suitable for both training and evaluation. We conduct a comprehensive evaluation with three types of models: zero-shot MLLMs, linters, and fine-tuned classifiers. We propose a new linter and a trained classifier that leverage the axis metadata extracted from a visualization by a fine-tuned DePlot model. Our results show that the task poses several challenges. MLLMs perform best on real-world visualizations, while fine-tuned classifiers have an edge on synthetic ones.
Limitations
We identify three main limitations in this work.
First, Misviz-synth is less diverse than its real-world counterpart, Misviz. It is limited to three chart types and does not include visualizations with multiple charts or multiple misleaders. Additionally, the use of the Matplotlib library imposes constraints, for instance, it does not support 3D pie charts. Nevertheless, the results from the fine-tuned classifiers indicate that the features learned from Misviz-synth transfer, to some extent, to the real-world visualizations in Misviz. Future work should explore alternative plotting libraries to broaden coverage across chart types and misleader categories, which should result in a better generalization to real-world visualizations.
Second, the set of 12 misleader categories addressed in this work represents only a subset of those found in practice. While our selection was carefully made based on three criteria, future efforts should expand the taxonomy to include more misleaders. Notably, it would be valuable to incorporate misleaders that require domain knowledge to detect, such as violating color conventions. Furthermore, our selection focuses on design misleaders. However, there are also reasoning misleaders, where no design rules are broken. Instead, they mislead through incomplete data or deceptive titles and annotations (Lisnic et al., 2023; Alexander et al., 2024). Such reasoning misleaders would require very different detection approaches and deserve more focus in future work.
Third, the boundaries between some misleader categories are not always clear-cut. For instance, inappropriate item order and inverted axis could be considered equivalent when a temporal axis is shown in reverse chronological order. This ambiguity is not captured by the current EM metric, which may slightly underestimate the true performance of models in detecting misleading visualizations.
Ethics statement
Social impact Misleading visualizations are frequently used by malicious actors to spread misinformation online, particularly on social media platforms (Correll and Heer, 2017; Lisnic et al., 2023). They may also result from unintentional design choices by chart designers. This work contributes both a dataset and new baselines to help mitigate the negative impact of misleading visualizations by enabling their automatic detection. The Misviz and Misviz-synth datasets are intended solely for academic research. As with other resources developed for misinformation detection, there is a potential risk of dual use, where malicious actors exploit detection models in adversarial settings to craft misleading content that evades detection. However, we believe that the potential benefits of this research, such as assisting chart designers and protecting readers, outweigh these risks.
Misinformation content Misviz includes real-world disinformation examples. To preserve the authenticity and diversity of misleading visualizations encountered in the wild, we did not filter or censor such content.
Dataset access Our code and dataset annotations are released under the Apache 2.0 and CC BY-SA 4.0 licenses, respectively. We do not hold the rights to the visualization images in Misviz. Therefore, we do not distribute them directly. Instead, we provide image URLs in the Misviz dataset file, along with a script to download them. To ensure long-term accessibility, we have verified that all images are available on the WaybackMachine and included the corresponding archive URLS in the dataset file.
AI assistants use AI assistants were used in this work to assist with writing by correcting grammar mistakes and typos.
Acknowledgments
This work has been funded by the LOEWE initiative (Hesse, Germany) within the emergenCITY center (Grant Number: LOEWE/1/12/519/03/05.001(0016)/72) and by the German Federal Ministry of Research, Technology and Space and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. Figures 2, 3, 13, and 14 have been designed using resources from Flaticon.com. We thank Liesbeth Allein, Luke Bates, Manisha Venkat, Max Glockner, and Vivek Gupta for our insightful discussions on misleading visualizations. We are grateful to Niklas Traser for preparing the annotation interface for crowdworkers. We also thank Shivam Sharma, Shivam Sharma (Junior), and Germán Ortiz for their feedback on an early draft of this work.
Appendix A Explanation of Misviz examples
Figure 4 shows the same misleading visualizations as in Figure 1, now overlaid with visual explanations highlighting the specific misleaders. We describe each case below:
(1) Misrepresentation In terms of percentages, Bernie Sanders is closer to Joe Biden than to Elizabeth Warren. However, the fourth bar (Sanders) is visually closer in height to the first (Warren) than the second (Biden), exaggerating the gap between Biden and Sanders.
(2) 3D The 3D effects make it difficult to visually compare values across years. For instance, it’s unclear which year had the highest number of singles sold between 1995 and 1997.
(3) Truncated axis The vertical axis begins at 36%, exaggerating the gap between pro- and anti-Brexit responses.
(4) Inappropriate use of pie chart The pie chart shows responses to three overlapping time categories: today, last year, and 1997. Because respondents may fall into more than one category, the percentages exceed 100%. The pie chart suggests the categories are distinct.
(5) Inconsistent binning size The two bars compare economic growth over different time spans. The first covers a longer period than the second, making the comparison misleading.
(6) Discretized continuous variable Germany and Denmark have similar values, but are shown in different colors. Meanwhile, Austria shares Germany’s color despite being further apart in value. The discrete color scale exaggerates differences between countries.
(7) Inconsistent tick intervals The dates on the horizontal axis are spaced unevenly. This distorts the slope of the trend line and gives a misleading impression of the average wedding costs progression over time.
(8) Dual axis The scores for John and Adam are plotted on separate vertical axes. While the visual gap on test 3 looks small on the left axis, the actual difference is 11%, not 1%.
(9) Inappropriate use of line chart A line chart is used to connect values across a categorical variable. The line implies a trend that does not exist.
(10) Inappropriate item order The dates on the horizontal axis are out of chronological order, which can mislead readers into thinking COVID-19 cases initially dropped and then rose. In fact, the reverse is true.
(11) Inverted axis The vertical axis increases from top to bottom. At first glance, it seems that enrollment in the Affordable Care Act is falling over time, while it is actually rising.
(12) Inappropriate axis range The vertical axis is so broad that the trend in global average temperature appears nearly flat. A more appropriate narrow axis range would reveal a clearer upward trend.
Appendix B Example of discarded visualizations
Figure 5: Examples of visualizations discarded from the corpus of Lo et al. (2022). Top: The misleading visualization is shown together with a correct pair for the purpose of educating readers about misleaders. Bottom: the visualization is of too low resolution, making the labels unreadable.
Figure 5 provides two examples from the corpus of Lo et al. (2022) that were not included in Misviz.
Appendix C Mapping of misleader categories between taxonomies
Table 5 presents the manual mapping applied between the misleader categories from the taxonomy of Lan and Liu (2025) to those of Lo et al. (2022).
Lan and Liu (2025) Lo et al. (2022)
Data-visual disproportion Misrepresentation
3D effect 3D
Truncated axis Truncated axis
Misuse of part-to-whole relationship (pie chart) Inappropriate use of pie chart
Uneven axis interval Inconsistent tick intervals
Dual axis Dual axis
Uneven data grouping Inconsistent binning size
Categorical encoding for continuous data Discretized continuous variable
Continuous encoding for categorical data (line chart) Inappropriate use of line chart
Inverted axis Inverted axis
Table 5: Mapping of misleader categories from the taxonomy of Lan and Liu (2025) to the taxonomy of Lo et al. (2022).
Appendix D Annotation interface and instructions
Figure 6: Main interface used by Prolific crowdworkers to assign misleader labels to a visualization.
Figure 6 shows the main interface used by Prolific crowdworkers to label the visualizations from r/dataisugly and r/dataisbeautiful.
Figures 7, 8, and 9 show the labeling instructions given to the crowdworkers.
Task 1: Chart types In this study, you will be asked to classify a chart into the correct categories (pie chart, line chart, bar chart, map, scatter plot, or other). You are expected to spend, on average, 10 to 15 seconds on each chart. Select one or more types of charts that appear in the image. You should always select at least one. You can choose multiple answers if the image contains more than one type.
Figure 7: Instructions for the first annotation task.
Task 2: Misleading visualizations In this study, your task is to detect misleaders present in a chart. Misleaders are design flaws that decrease the ability of readers to interpret charts correctly. You are expected to spend, on average, 30 seconds to 1 minute on each chart. Misleaders are chart design elements that can lead readers to interpret a wrong message that does not correspond to the underlying data. This study covers 12 types of misleaders. Each chart contains one to three misleaders (one being the most frequent case). Your task is to select the correct misleader(s) for each chart. We present the types of misleaders below. Important remarks - To perform this task, you need to pay attention to all aspects of the chart: axes and their labels, titles, legends, and proportions. Misleaders often reside in small details. -It is essential to select the correct misleader(s) that apply to the image Most charts contain only one misleader. In some cases, there will be two or three.
Figure 8: Instructions for the second annotation task.
Task 3: Non-misleading visualizations In this study, your task is to verify that a chart is not misleading, i.e., it does not have significant design flaws that decrease the ability of readers to interpret charts correctly. Your participation will help us build AI systems capable of classifying charts as misleading or not. You are expected to spend, on average, 30 seconds to 1 minute on each chart. You will have to review charts that come from an online community. The majority of the charts are expected to be of high quality. However, some of them might be misleading. Misleaders are chart design elements that can lead readers to interpret a wrong message that does not correspond to the underlying data. There are 12 types of misleaders to consider, and they are explained below. If the chart is not misleading (i.e., in MOST CASES), you should leave the answer field blank. If the chart is misleading, select the misleaders that apply (up to three misleaders). Important remark - To perform this task, you need to pay attention to all aspects of the chart: axes and their labels, titles, legends, and proportions. Misleaders often reside in small details. - You should not select a misleader for each chart! Most of them are not misleading. You should only detect the few that are misleading!
Figure 9: Instructions for the third annotation task.
Figure 10: Four visualizations from Misviz. Top: the embedded text is not written in English. Bottom: the visualization is not a screenshot but a photograph taken of an electronic device or printed material.
Figure 11: Two visualizations from Misviz-synth with their corresponding axis metadata.
Appendix E Detailed dataset statistics
Tables 6 and 7 report the distributions of misleader categories and chart types in Misviz and Misviz-synth, respectively. The percentages do not sum to 100% for Misviz as a visualization may contain more than one misleader category or chart type.
Of the 1,791 misleading visualizations in Misviz, 353 originate from the corpus of Lo et al. (2022), 964 were collected from the WTF Visualizations website (Lan and Liu, 2025), and the remaining 474 were sourced from r/dataisugly and annotated by crowdworkers.
Misleader category Misviz Misviz-synth
Misrepresentation 32.6 8.5
3D 14.0 4.1
Truncated axis 8.7 2.0
Inappropriate use of pie chart 6.2 2.3
Inconsistent tick intervals 5.1 24.5
Dual axis 3.0 1.5
Inconsistent binning size 2.8 4.5
Discretized continuous variable 1.9 0.0
Inappropriate use of line chart 1.8 2.7
Inappropriate item order 1.5 7.7
Inverted axis 1.2 18.8
Inappropriate axis range 1.1 1.3
Table 6: Percentage of instances containing a given misleader category.
Chart type Misviz Misviz-synth
Bar 37.0 46.1
Line 20.0 46.7
Pie 23.4 7.2
Map 4.9 0.0
Scatter plot 4.3 0.0
Other 16.5 0.0
Table 7: Percentage of instances containing a given chart type.
Appendix F Examples of language and image type diversity in Misviz
Figure 10 shows four instances highlighting the diversity of Misviz. The two visualizations at the top have embedded text written in German and Estonian. The bottom visualizations are photographs of a visualization shown on television screen and on a printed document.
Appendix G Misviz-synth column types
Table 8 lists the 10 column types used in the first step of the Misviz-synth synthetic visualization generation process. Each column is assigned exactly one primary type, which determines its core data role. Additionally, columns can be assigned one or more secondary types, which describe extra attributes that help determine appropriate chart types and potential misleaders.
Column type Characteristic
Primary column types
Temporal Contains temporal values
Categorical Contains categorical values
Numerical Contains numerical values
Secondary column types
Datetime Datetime objects
Evenly spaced, unique temporal Evenly spaced temporal values
Country Country names
Unique object Non-numerical values appear only once
Is part of whole Numerical values sum to 1 or 100
Numerical percentage Explicit percentage values
(e.g., marked with %)
Potential percentage Potential percentage values
(e.g., value range is [0,1] or [0,100])
Table 8: Primary and secondary column types used during synthetic visualization generation
Appendix H Examples of visualizations with axis metadata
Figure 11 provides two examples of visualizations from Misviz-synth with the corresponding axis metadata. The latter is represented as a table and stored as a JSON file with column names as keys.
Task prompt for misleader detection with zero-shot MLLMs You are an expert in data visualization analysis. Your task is to identify misleaders present in the given visualization. Please carefully examine the visualization and detect its misleaders. Provide all relevant misleaders, up to three, as a comma separated list. In most cases only one misleader is relevant. If you detect none of the above types of misleaders in the visualization, respond with “no misleader”. The available misleaders to select are, by alphabetical order: - discretized continuous variable: a map displays a continuous variable transformed into a categorical variable by cutting it into discrete categories, thus exaggerating the difference between boundary cases. - dual axis: there are two independent y-axis, one on the left and one on the right, with different scales. - inappropriate axis range: the axis range is too broad or too narrow. - inappropriate item order: instances of a variable along an axis are in an unconventional, non-linear or non-chronological order. - inappropriate use of line chart: a line chart is used in inappropriate or unconventional ways, e.g., using a line chart with categorical variables, or encoding the time dimension on the y-axis. - inappropriate use of pie chart: a pie chart does not display data in a part-to-whole relationship, e.g., its shares do not sum to 100%. - inconsistent binning size: a variable, such as years or ages, is grouped in unevenly sized bins. - inconsistent tick intervals: the ticks values in one axis are evenly spaced but their values are not, e.g., the tick value sequence is 10, 20, 40, 45. - inverted axis: an axis is displayed in a direction opposite to conventions, e.g., the y-axis displays values increasing from top to bottom or the x-axis displays values increasing from right to left. - misrepresentation: the value labels displayed do not match the size of their visual encodings, e.g., bars may be drawn disproportionate to the corresponding numerical value. - truncated axis: an axis does not start from zero, resulting in a visual exaggeration of changes in the dependent variable with respect to the independent variable. - 3d: the visualization includes three-dimensional effects. Provide only the final answer, without additional explanation.
Figure 12: Task prompt for zero-shot misleader detection with zero-shot MLLMs.
Appendix I List of random design variations
All visualizations: color variation of the background, variations of the title template, variation of font type and size, and variation of chart size.
Bar and line charts: addition of minor ticks in addition to major ones, positioning of the vertical axis (left or right), variation in tick shapes, adding value labels on top of bars, variation of the tick step size, addition of chart borders, andaddition of horizontal grid lines.
Bar charts only: sorting bars by values or by category name, hiding the vertical axis, placing labels on top of the bars or within the bars, using horizontal or vertical value labels, variation of bar colors.
Line charts only: filling area below the line with a color, variation in line style and color, and addition of horizontal or vertical grid lines.
Pie charts only: placing data label next to the pie slices or placing them or in a legend.
Appendix J Zero-shot prompts
Figure 12 shows the task prompt provided to MLLMs to detect misleading visualizations. The prompt starts by providing task instructions, followed by the definitions of the misleaders in alphabetical order. For Misviz-synth, the misleader category discretized continuous variable is removed from the prompt.
Appendix K Linter rules
We explain below the linter rules to detect some of the misleaders in Misviz and Misviz-synth. These rules are implemented in Python and were designed on the validation set of Misviz-synth to maximize precision over recall.
Truncated axis A visualization has a truncated axis if one or more of its sorted vertical axes starts with a numerical value strictly higher than 0.
Inverted axis A visualization has an inverted axis if one or more of its sorted axes is the reverse of the default axis order.
Dual axis A visualization has a dual axes if it has two vertical axes with different axis tick labels.
Inappropriate item order A visualization has an inappropriate item order if it has a temporal axis with dates shuffled in random order. This covers only a specific case of inappropriate item order.
Inconsistent tick intervals A visualization has inconsistent tick intervals if the distance between the ticks on one or more of its numerical or temporal axes is not constant.
Inconsistent binning size A visualization suffers from inconsistent binning size if one axis shows categorical bins of unequal size.
Misviz Misviz-synth
Acc Pre Rec F1 EM PM Acc Pre Rec F1 EM PM
Image 69.6 $\pm$ 0.9 72.2 $\pm$ 0.2 94.3 $\pm$ 1.3 45.6 $\pm$ 0.4 27.1 $\pm$ 1.9 34.0 $\pm$ 2.3 79.2 $\pm$ 0.2 83.8 $\pm$ 0.4 90.9 $\pm$ 1.0 66.0 $\pm$ 0.7 76.8 $\pm$ 0.1 76.8 $\pm$ 0.1
Image w. axis (DePlot[Uncaptioned image]) 64.3 $\pm$ 3.0 71.4 $\pm$ 0.8 84.4 $\pm$ 4.7 46.7 $\pm$ 1.3 22.1 $\pm$ 1.1 27.7 $\pm$ 1.4 81.6 $\pm$ 0.5 86.6 $\pm$ 0.5 90.4 $\pm$ 1.3 71.7 $\pm$ 0.5 80.4 $\pm$ 0.7 80.4 $\pm$ 0.7
Table 9: Average performance with standard deviations (%) of the fine-tuned classifiers on the test sets of Misviz and Misviz-synth for instances with bar, line, or pie charts.
Figure 13: Error examples on the test set of Misviz. The predictions made by GPT-4.1 are shown in blue while the ground truth is shown in green.
Appendix L Results with standard deviations of fine-tuned classifiers
Table 9 reports the average results with standard deviations of the fine-tuned classifiers on the subset of Misviz with bar, pie, and line charts and on Misviz-synth.
Figure 14: Error examples on the test set of Misviz-synth. The predictions made by GPT-4.1 are shown in blue while the ground truth is shown in green.
Appendix M Additional implementation details
Training sets To train the DePlot axis extractor and the MLP classification head, we split the Misviz-synth train set into two parts. The larger portion, representing 80% of the data, is used to fine-tune DePlot. The remaining 20% is used to train the MLP classification head. To increase DePlot’s sensitivity to varying tick intervals and step sizes, we add 3,500 additional synthetic visualizations to the axis extraction train set. These are specifically designed to introduce diverse axis spacings.
DePlot fine-tuning We fine-tune DePlot for axis metadata extraction using two H110 GPUs. We retain the original DePlot architecture and loss function (Liu et al., 2023). The model is trained for 4 epochs using the Adam optimizer (Kingma and Ba, 2015), a learning rate of 5e-5, and a batch size of 4. To support memory-efficient training, we apply LoRA (Hu et al., 2022) on the query and value projection layers. To enhance generalization, we apply data augmentation through random rotations and perspective transformations.
Fine-tuned classifiers The MLP classification head is trained for up to 300 epochs with early stopping. We use a batch size of 256, a learning rate of 5e-5, and the Adam optimizer. To account for label imbalance, we apply a weighted loss based on the frequency of each misleader. The MLP has a hidden dimension of 1,024.
Appendix N Misviz results per data source
Corpus by WTF Visualization
Lo et al. (2022) (Lan and Liu, 2025) r/dataisugly
Rec EM PM Rec EM PM Rec EM PM
Qwen2.5-VL-7B 47.1 7.6 17.9 24.7 2.7 3.4 31.8 7.2 7.2
Qwen2.5-VL-32B 96.0 9.2 12.6 96.9 2.5 3.3 93.5 2.7 2.7
Qwen2.5-VL-72B 93.7 31.5 55.5 82.8 20.4 21.5 83.2 32.5 32.5
InternVL3-8B 92.6 18.1 37.2 82.6 22.0 22.7 87.0 29.8 29.8
InternVL3-38B 71.8 23.7 49.2 53.9 34.6 35.5 51.7 21.9 21.9
InternVL3-78B 80.7 25.4 46.0 62.2 23.7 24.3 59.6 21.6 21.6
Gemini-1.5-Flash 74.6 27.1 48.1 44.9 19.5 20.5 57.9 18.5 18.5
Gemini-1.5-Pro 83.4 35.5 51.7 81.0 37.6 38.6 64.4 28.4 28.4
GPT-4.1 96.2 43.5 73.1 94.2 58.8 60.2 91.4 57.9 57.9
GPT-o3 93.1 50.4 75.2 88.9 63.2 64.2 88.0 62.7 62.7
Table 10: Zero-shot MLLMs performance (%) on the test set of Misviz, separated by the data source of the misleading visualizations. The data source with the lowest EM for each MLLM is marked in bold.
Table 10 reports the performance of zero-shot MLLMs across three subsets of the Misviz test set, corresponding to its three sources of misleading visualizations. Performance varies significantly across subsets, but no single source is consistently easier or harder across all models. This highlights the complementary value of each data source to the overall Misviz benchmark. Interestingly, the subset with the lowest EM is often consistent within the same model family, suggesting shared weaknesses in handling visualizations from the same data source.
Appendix O Error examples
Figures 13 and 14 show example of errors made by GPT-4.1 on the test sets of Misviz and Misviz-synth, respectively. One example is shown for each error category. Each example shows the prediction of GPT-4.1 in blue, the correct answer in green, and the error category in gray.